Can you help me solve this application? I'm trying to solve applications involving parabolas.
The length of a rectangle is twice the width. The area is 16m^2. Find the length to the nearest hundredth of a meter.
Thanks
Liz

Answer by ankor@dixie-net.com(22740)

The length of a rectangle is twice the width. The area is 16 m^2. Find the length to the nearest hundredth of a meter.
:
Let x = the width
Then 2x = the length
:
Area:
x*2x = 16
2x^2 = 16
x^2 = 16/2
x^2 = 8
x = SqRt(8)
x = 2.828
:
Length = 2* 2.828 = 5.66 meters
:
Check 2.83*5.66 = 16.02 because of rounding off